OpenStudy (anonymous):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2, find y2. (18, 2) and (x2, y2), if x2=y2 a.y2 = +-6 c.y2 = 0 b.y2 = +-1 d.y2 = -6

OpenStudy (sleepyhead314):

if x2 = y2 then x1y1=y2y2 so 18*2 = y2y2 solve for y2

OpenStudy (anonymous):

explain

OpenStudy (sleepyhead314):

(y2)^2 = 18*2 so y2 = square root of (18 times 2) = ?

OpenStudy (anonymous):

so +-6

OpenStudy (sleepyhead314):

correct

OpenStudy (anonymous):

can you help me with the next one please

OpenStudy (sleepyhead314):

is this a test or something?

OpenStudy (anonymous):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (x, 25) and (5, 125) a.x = +-16 c.x = 17 b.x =+-25 d.x = 25

OpenStudy (anonymous):

no homework

OpenStudy (sleepyhead314):

plug in x1=x y1= 25 x2= 5 y2=125 into x1*y1 = x2*y2

OpenStudy (anonymous):

x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (3, y) and (18, 6) a.y = 36 c.y = 32 b.y = -42 d. y = -16

OpenStudy (sleepyhead314):

one question at a time please I really don't like it when people aim to only get answers and refuse to learn when I am trying to teach.

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

the answer is 1

OpenStudy (sleepyhead314):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (x, 25) and (5, 125) plug in x1=x y1= 25 x2= 5 y2=125 into x1*y1 = x2*y2 what would you get?

OpenStudy (anonymous):

25

OpenStudy (sleepyhead314):

yes

OpenStudy (anonymous):

would it just be 25 or +-25

OpenStudy (sleepyhead314):

there were no square roots involved when going through the process so only positive

OpenStudy (anonymous):

next question

OpenStudy (sleepyhead314):

x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (3, y) and (18, 6) plug in x1=3 y1=y x2=18 y2=6 into x1*y1 = x2 *y2 what you do get?

OpenStudy (anonymous):

a 36

OpenStudy (sleepyhead314):

yes

OpenStudy (sleepyhead314):

are you getting the hang of this? do you see the pattern of how to solve these sorts of problems now?

OpenStudy (anonymous):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (3.6, 5) and (7.2, y)

OpenStudy (sleepyhead314):

see how (x1, y1) and (x2, y2) match up with (3.6, 5) and (7.2, y) ^ ^ ^ ^ plug it in that way into x1*y1 = x2*y2

OpenStudy (anonymous):

2.5

OpenStudy (sleepyhead314):

correct

OpenStudy (anonymous):

z varies jointly as x and y and inversely as w. Write the appropriate combined-variation equation, and find z for the given values of x, y and w. z = 10 when x = 5, y = -2 and w = 3; x = 8, y = 6 and w = -12

OpenStudy (sleepyhead314):

eh sorry but you're better off posting that in a separate question post I am not familiar with inversely and jointly stuff, sorry

OpenStudy (anonymous):

ok